OpenStudy (anonymous):

Factor 6x^4+15x^3-9x^2

OpenStudy (mathmate):

Factor out the common factor 3x^2 and you're left with a quadratic which you probably know how to factor.

OpenStudy (anonymous):

Thanks

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

Factor 6x^4+15x^3-9x^2 3x^2(2x^4+5x-3) x(3x+6)1(x-1) ?? can someone tell me what im doing wrong apparently 2x^4 is supposed to be 2x^2 and i dont know why as well

myininaya (myininaya):

so if you factor out 3x^2 divide each term inside the parenthesis by 3x^2 like so \[3x^2(\frac{6x^4}{3x^2}+\frac{15x^3}{3x^2}-\frac{9x^2}{3x^2})\] simplify each of the fractions there

OpenStudy (anonymous):

oh ok thanks, ill try it out in a bit

OpenStudy (anonymous):

i think im mixing concepts together from all the vids ive watched

OpenStudy (mathmate):

Did you get it yet? @5454354366464

OpenStudy (anonymous):

no :( after 3x^2(2x^2+5x-3) dont u 2(-3) for product then keep the 5x as is product = -6 and sum = 5

OpenStudy (anonymous):

so it would be 6 and -1 for something that multiples and adds

OpenStudy (anonymous):

(2x+6)(2x-1) 3x^2(x+3)(2x-1) ok i got it thanks :D

OpenStudy (mathmate):

Yep, you got it! Good job! :)